hi i need help for this equation and how to to arrive at the answer.
solve (find values) for A, B, C, D, F
(x^4 - 4x^2 - 7x - 28) = A(x^B + C)(x^D - F)

Answer by khwang(438)

x^4 - 4x^2 - 7x - 28
= x^2(x^2-4) - 7(x+4)
= x^2(x-2)(x+2) - 7(x+4) =????
Also try synthnic division, there is no factor of deg 1
and clearly it is not a product of two 2nd degree rational polynomials.
Plz check your left hand side.
Kenny